OpenStudy (anonymous):

Will the function \[\frac{ (x+1)(x-2) }{ (x-3) }\], have the slant asymptote???

Parth (parthkohli):

Yeah, at \(y = x + 2\)

OpenStudy (anonymous):

(x+1)(x-2)=x^2-x-2

OpenStudy (anonymous):

thought so.

OpenStudy (anonymous):

Now, divide x^2-x-2 by x-3

OpenStudy (anonymous):

thanks for all responses

Parth (parthkohli):

Remember: the quotient is always our slant asymptote.

Parth (parthkohli):

Sorry... it'd actually be at y = x + 6

OpenStudy (anonymous):

the reason I asked is because my texts solution only shows an asymptote at x=3

Parth (parthkohli):

Ok:)

OpenStudy (anonymous):

so my text is wrong?